# RefAM: Attention Magnets for Zero-Shot Referral Segmentation

### v1

@@ 4.1 SOTA Comparison
AM NP SB RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
[…]
- 33.89 44.66 34.14 35.12 37.69 33.75 42.93 42.44
- 37.60 41.81 34.22 38.53 42.60 35.66 42.59 42.75
- - 32.99 35.22 31.98 34.47 37.05 33.62 41.83 41.00
- - 35.54 39.80 32.86 37.12 40.89 34.32 38.66 40.90
- - - 29.14 31.49 29.21 31.61 34.29 30.65 34.09 35.81
4.2 Ablations
In Tabs. 5 and 5, we decouple $\mathcal{J}\&\mathcal{F}$ mask evaluation from our predicted points by introducing the point accuracy (PA) metric, which considers a point correct if it falls within the ground-truth mask.
Influence of Attention Magnets. Including stop words in attention map aggregation results in overly diffuse localization. As shown in Tabs. 5 and 6, introducing and then filtering our attention magnets (AM) out from the referring expressions improves predicted point accuracy from 59.9 to 68.9 and raises the $\mathcal{J}\&\mathcal{F}$ metric by 3.2 points on RVOS. Moreover, we observe consistent gains across settings when attention magnets are appended. Fig. 5 further illustrates how redistributing background activations followed by filtering, produces sharper and more focused attention maps.
Noun Prase and Spatial Bias. We conduct an ablation study to disentangle the contributions of spatial bias and noun phrase encoding, as shown in  Tab. 6 and Tab. 5. We use the same preprocessing strategy as HybridGL (Liu & Li, 2025) to extract noun phrases and spatial relations from the referring expression, ensuring a fair comparison. When combined, the two components with our attention magnets yield the best performance across all benchmarks, confirming their complementary roles in grounding referring expressions. See Appendix C for more details.
@@ 5 Conclusion
Use of Large Language Models (LLMs)
We used an LLM as an auxiliary tool in two limited capacities. For dataset preparation, the model generated captions for DAVIS, Ref-YouTube-VOS, and MeViS test videos, as well as semantically related prompt variations. All generated outputs were carefully screened and curated by the authors before use. For writing, the authors drafted all sections of the paper, and the LLM was employed only for copy-editing and stylistic refinement. The model was never used to introduce technical content, conduct experiments, or contribute to the research methodology.
@@ Appendix A Additional Experiments
A.1 Feature Collection with Inversion
In the main paper, we extract diffusion features without applying inversion. For completeness, we also report results obtained when performing inversion prior to feature collection.
Setup. All experiments were conducted on a single NVIDIA A100 GPU. The peak memory usage was approximately 37GB for both Flux and Mochi models. For RefAM + inversion, the inversion step requires 6–9 seconds per input, while subsequent denoising (28 steps) requires 16–20 seconds. Video segmentation with 50 frames using SAM2 adds 3–4 seconds. We use a batch size of 1 for the visual input, but our framework supports multiple referring expressions per image or video. In practice, we process 5–8 expressions in parallel without a hard limit. Overall, feature extraction with RefAM + inversion is roughly 60% faster than the commonly used SD+DINO pipeline.
Feature collection via inversion. Generative diffusion models are trained to synthesize images from pure noise, and directly perturbing a real image with Gaussian noise does not guarantee that the resulting latent follows a trajectory observed during generation. This mismatch often causes reconstructions that diverge from the input in content or style (Hertz et al., 2023; Rout et al., 2025). To address this, we adopt rectified flow inversion (Rout et al., 2025), which introduces a controlled backward process ensuring trajectory consistency. The latent dynamics are defined as
$\frac{dY_{t}}{dt}=-v_{1-t}(Y_{t})+\gamma\big(u_{t}(Y_{t}\mid y_{1})+v_{1-t}(Y_{t})\big),$
where $u_{t}(\cdot)$ is the forward vector field, $v_{1-t}(\cdot)$ its reverse, and $\gamma\in[0,1]$ a correction factor. This inversion yields structured noise latents $X_{T}$ that, when denoised, reconstruct the original input $X_{0}$ while preserving semantic fidelity.
Procedure. Our feature collection consists of two stages: (i) invert the clean input $X_{0}$ into a structured noise latent $X_{T}$ via rectified flow inversion; (ii) apply standard rectified flow denoising to $X_{T}$, collecting cross-attention maps from intermediate DiT blocks. To further improve semantic quality, we generate a caption for each input and condition both inversion and denoising on this caption. This ensures that the extracted features encode meaningful semantics aligned with downstream referring segmentation tasks.
Results and trade-off. We compare feature extraction with and without inversion on standard referring expression benchmarks in Table 7. While both variants establish new state-of-the-art zero-shot results, inversion consistently improves performance across RefCOCO, RefCOCO+, and RefCOCOg, particularly in terms of mIoU. However, these gains come at the cost of additional latency: inversion adds 6–9 seconds per input before denoising. Thus, practitioners may choose between the non-inversion variant for efficiency, or the inversion-based variant for maximal accuracy, depending on application requirements.
Metric Method Vision Backbone Pre-trained Model RefCOCO RefCOCO+ RefCOCOg
@@ A.1 Feature Collection with Inversion
oIoU RefAM (ours) DiT SAM, FLUX 46.91 52.30 43.88 38.57 42.66 34.90 45.53 44.45
RefAM + Inversion (ours) DiT SAM, FLUX 47.99 52.22 43.99 40.29 46.05 36.71 47.08 47.38
mIoU RefAM (ours) DiT SAM, FLUX 57.24 59.78 53.32 43.59 47.28 38.77 47.11 48.35
RefAM + Inversion (ours) DiT SAM, FLUX 59.12 60.21 55.02 46.86 51.13 42.30 49.42 50.55
Table 7: Zero-shot referring segmentation results on RefCOCO, RefCOCO+, and RefCOCOg. We compare our method with and without inversion. Both variants achieve state-of-the-art performance, and inversion further improves consistency and semantic alignment, particularly in mIoU. Best and second-best results are shown in bold and underlined, respectively.
@@ A.3 Comparison to ConceptAttention on Image Segmentation Task
GradCAM (Selvaraju et al., 2017b) CLIP ViT 64.44 40.82 71.60 70.44 44.90 76.80
@@
DAAM (Tang et al., 2023b) SDXL UNet 78.47 64.56 88.79 72.76 55.95 88.34
DAAM (Tang et al., 2023b) SD2 UNet 64.52 47.62 78.01 64.28 45.01 83.04
@@
[…]
Table 11: Our method, RefAM with inversion, consistently outperforms a range of interpretability techniques based on Diffusion, DINO, CLIP ViT, and Flux DiT on both ImageNet-Segmentation and PascalVOC (Single Class). The performance numbers for the other methods are taken directly from ConceptAttention, and we follow the same evaluation procedure to ensure fair comparison.
Appendix B Stop Word Filtering & Additional Stop Words
@@
$\hat{H}_{k}=\text{softmax}_{k}(H_{k}).$
This normalization implies that for each visual patch we define a probability distribution that associate it with the token having the highest softmax score relative to that patch. Given that the referral expression corresponds specifically to a particular region or element within the visual input, it follows that visual areas not directly associated with the referral expression must be attributed to other tokens. We observe that words with minimal semantic significance, such as stop words, often represent the broader context or background elements of the scene relative to the specific referral expression.
@@
i; me; my; myself; we; our; ours; ourselves; you; your; yours; yourself; yourselves; he; him; his; himself; she; her; hers; herself; it; its; itself; they; them; their; theirs; themselves; what; which; who; whom; this; that; these; those; am; is; are; was; were; be; been; being; have; has; had; having; do; does; did; doing; a; an; the; and; but; if; or; because; as; until; while; of; at; by; for; with; about; against; between; into; through; during; before; after; above; below; to; from; up; down; in; out; on; off; over; under; again; further; then; once; here; there; when; where; why; how; all; any; both; each; few; more; most; other; some; such; no; nor; not; only; own; same; so; than; too; very; s; t; can; will; just; don; should; now; ,; .; _;
Appendix C Details of Noun Phrase and Spatial Bias Extraction
We adopt the same preprocessing strategy as HybridGL (Liu & Li, 2025) to extract noun phrases and spatial cues from referring expressions. Following their approach, we parse each input sentence into object-centric noun phrases (e.g., “the man”, “a red car”) and spatial relations (e.g., “left of”, “behind”, “top”). The noun phrases are then encoded with the text encoder and compared against diffusion-derived visual features, guiding attention toward semantically relevant regions.
Spatial cues are incorporated as lightweight spatial priors. Relative relations (“left of the dog”) are modeled by comparing bounding box centroids of candidate regions, while absolute terms (“top left”, “bottom right”) are mapped to normalized positional masks over the image grid. This procedure mirrors the spatial relationship guidance in HybridGL but operates directly on cross-attention features extracted from diffusion transformers.
As shown in our ablation study (Table 6), both components contribute complementary gains. Spatial bias alone improves localization accuracy, while noun phrase extraction enhances semantic alignment. When combined with our attention magnet mechanism, the two yield the strongest results across benchmarks. We further confirm this effect in video segmentation benchmarks (Table 5), where the same preprocessing consistently benefits temporal grounding.
@@ Appendix F Qualitative Examples
Figure 10: Qualitative examples. VROS task, evaluated on Ref-DAVIS17. From left to right: first frame of the video with the corresponding ref.expression on the top, avg. attention map from our RefAM + inversion features, segmentation outputs with SAM2.

### v2

@@ 4.1 Quantitative Analysis
AM NP SB RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
[…]
- 39.99 41.32 39.11 32.56 35.54 29.22 38.57 40.11
- 31.89 35.82 29.17 34.29 38.33 29.82 37.48 38.87
- - 28.45 31.61 27.76 32.22 35.09 28.96 36.72 38.23
- - 31.13 34.31 28.78 33.19 36.03 29.16 32.39 32.11
- - - 27.45 30.12 27.48 30.96 32.74 28.30 31.55 31.91
4.2 Ablations
In Tabs. 5 and 6, we decouple $\mathcal{J}\&\mathcal{F}$ mask evaluation from our predicted points by introducing the point accuracy (PA) metric, which considers a point correct if it falls within the ground-truth mask.
Influence of Attention Magnets. Including stop words in attention map aggregation results in overly diffuse localization. As shown in Tabs. 5 and 4, introducing and then filtering our attention magnets (AM) out from the referring expressions improves predicted point accuracy from 59.9 to 68.9 and raises the $\mathcal{J}\&\mathcal{F}$ metric by 3.2 points on RVOS. Moreover, we observe consistent gains across settings when attention magnets are appended. Fig. 5 further illustrates how redistributing background activations followed by filtering, produces sharper and more focused attention maps.
@@ 4.2 Ablations
Noun Phrase and Spatial Bias. We conduct an ablation study to disentangle the contributions of spatial bias and noun phrase encoding, as shown in Tabs. 5 and 4. To extract noun phrases and spatial relations from the referring expression, we utilize the spaCy library. When combined, the two components with our attention magnets yield the best performance across all benchmarks, confirming their complementary roles in grounding referring expressions. See Appendix C for more details.
@@
4.3 Generalization & Backbone Analysis
To verify that our performance gains stem from the proposed methodology rather than solely from the specific FLUX backbone, we extend our evaluation to Stable Diffusion 3.5 (SD3.5) which employs a hybrid text encoding scheme utilizing both CLIP and T5 encoders, allowing us to decouple their contributions and analyze the source of semantic grounding.
T5 vs. CLIP Encoders. As shown in Table 7, utilizing the T5 encoder alone yields significantly better performance than CLIP. We observe that T5 is structure-aware: removing stop words (w/o RefAM) causes a sharp performance drop (e.g., -5.2% mIoU on RefCOCO TestA). Conversely, CLIP acts effectively as a “bag-of-words” model (Yuksekgonul et al., 2023); removing stop words often improves its performance, indicating it fails to utilize syntactic structure for fine-grained grounding. This validates our design choice: RefAM exploits the fine-grained structural alignment present in modern T5-based DiTs, which is largely absent in CLIP-based dual-encoders. We provide the complete evaluation across all datasets in Sec. A.1.
T5 Encoder CLIP Encoder Combined
Metric w/ RAM w/o RAM w/ RAM w/o RAM w/ RAM w/o RAM
oIoU 41.5 36.3 38.3 36.9 44.8 39.1
mIoU 49.7 43.9 46.3 44.8 52.6 46.9
Table 7: Backbone Analysis on SD3.5 (RefCOCO TestA). T5 provides structural understanding (sensitive to Stop Words), while CLIP behaves like a Bag-of-Words. RAM denotes RefAM.
@@ Appendix A Additional Experiments
A.1 Analysis on Stable Diffusion 3.5
In Sec. 4.3, we present the ablation study on the RefCOCO TestA split to demonstrate the structural differences between T5 and CLIP encoders. In Table 8, we provide the complete evaluation across all splits of RefCOCO, RefCOCO+, and RefCOCOg.
The results consistently confirm our findings: the T5 encoder is structure-aware and suffers significant performance drops when stop words are removed (w/o RefAM). In contrast, the CLIP encoder acts largely as a “bag-of-words” model, often showing insensitivity or even slight improvements when stop words are removed, but failing to achieve the peak performance of the T5 encoder on complex splits.
T5 CLIP AM RefCOCO RefCOCO+ RefCOCOg
@@
- 39.39 41.51 36.46 32.29 36.51 28.23 37.04 38.29
- - 35.34 36.29 34.98 30.01 32.46 26.74 35.67 37.72
- 36.41 38.34 34.46 31.97 35.73 27.93 34.91 37.39
- - 35.51 36.87 34.28 31.03 34.86 27.56 34.80 36.45
41.93 44.80 38.37 33.40 38.03 28.90 37.38 39.31
- 38.00 39.07 36.94 32.14 35.71 28.10 36.39 39.01
Table 8: Full ablation of Text Encoders in SD3.5 across RefCOCO, RefCOCO+, and RefCOCOg (Metric: oIoU). T5 consistently outperforms CLIP and is sensitive to RefAM (Stop Words), confirming it drives structural grounding.
@@ A.3 Comparison to ConceptAttention on Image Segmentation Task
GradCAM (Selvaraju et al., 2020) CLIP ViT 64.44 40.82 71.60 70.44 44.90 76.80
@@
DAAM (Tang et al., 2022) SDXL UNet 78.47 64.56 88.79 72.76 55.95 88.34
DAAM (Tang et al., 2022) SD2 UNet 64.52 47.62 78.01 64.28 45.01 83.04
@@
Table 12: Our method consistently outperforms a range of interpretability techniques based on Diffusion, DINO, CLIP ViT, and Flux DiT on both ImageNet-Segmentation and PascalVOC (Single Class). The performance numbers for the other methods are taken directly from ConceptAttention, and we follow the same evaluation procedure to ensure fair comparison.
Appendix B Additional Stop Words & Filtering
@@
^H_k = softmax_k(H_k).
This normalization implies that for each visual patch we define a probability distribution that associates it with the token having the highest softmax score relative to that patch. Given that the referral expression corresponds specifically to a particular region or element within the visual input, it follows that visual areas not directly associated with the referral expression must be attributed to other tokens. We observe that words with minimal semantic significance, such as stop words, often represent the broader context or background elements of the scene relative to the specific referral expression.
@@
i; me; my; myself; we; our; ours; ourselves; you; your; yours; yourself; yourselves; he; him; his; himself; she; her; hers; herself; it; its; itself; they; them; their; theirs; themselves; what; which; who; whom; this; that; these; those; am; is; are; was; were; be; been; being; have; has; had; having; do; does; did; doing; a; an; the; and; but; if; or; because; as; until; while; of; at; by; for; with; about; against; between; into; through; during; before; after; above; below; to; from; up; down; in; out; on; off; over; under; again; further; then; once; here; there; when; where; why; how; all; any; both; each; few; more; most; other; some; such; no; nor; not; only; own; same; so; than; too; very; s; t; can; will; just; don; should; now; ,; .; _
Appendix C Noun Phrase and Spatial Bias Extraction
We adopt a preprocessing strategy using the spaCy library to extract noun phrases and spatial cues from referring expressions. Specifically, we parse each input sentence into object-centric noun phrases (e.g., “the man”, “a red car”) and spatial relations (e.g., “left of”, “behind”, “top”). The noun phrases are then encoded with the text encoder and compared against diffusion-derived visual features, guiding attention toward semantically relevant regions.
Spatial cues are incorporated as lightweight spatial priors. Relative relations (“left of the dog”) are modeled by comparing bounding box centroids of candidate regions, while absolute terms (“top left”, “bottom right”) are mapped to normalized positional masks over the image grid.
As shown in our ablation study (Tab 4), both components contribute complementary gains. Spatial bias alone improves localization accuracy, while noun phrase extraction enhances semantic alignment. When combined with our attention magnet mechanism, the two yield the strongest results across benchmarks. We further confirm this effect in video segmentation benchmarks (Table 5), where the same preprocessing consistently benefits temporal grounding.
@@ Appendix F Qualitative Examples
Figure 10: Qualitative examples. VROS task, evaluated on Ref-DAVIS17. From left to right: first frame of the video with the corresponding ref.expression on the top, avg. attention map from RefAM features, segmentation outputs with SAM2.